I need help with this question, Please
The height of a building is 1651 ft. how long would it take an object to fall to the ground from the top. Use s = 16t^2, where s is the distance, in feet, traveled by an object falling freely from rest in t seconds.
Thank you

Answer by richwmiller(17219)

1651=16t^2
1651/16=t^2
103.188=t^2
10.16=t